# An approximate Kappa generator for particle simulations

### v1

@@ I Introduction
In space physics, the Kappa distribution is arguably the most important velocity distribution after the the Maxwell distribution. Since it was empirically introduced in space during the 1960s, (Olbert, 1968; Vasyliunas, 1968) the Kappa distributions have been extensively observed throughout the heliosphere (Pierrard & Lazar, 2010; Livadiotis, 2017) and they have become increasingly popular. Theoretically, it has also been established that the Kappa distribution is related to Tsallis statistics.(Livadiotis & McComas, 2009) Since the Kappa distribution maximizes non-extensive entropy, it may represent a terminal state after plasma long-range interactions.(Tsallis, 2023) Driven by spacecraft observations and various theoretical motivations, the kinetic modeling of Kappa-distributed plasmas is growing in its importance.(Park et al., 2013; Gedalin & Ganushkina, 2022; Fitzmaurice, 2024; Nechaev et al., 2025)
@@
In kinetic simulations, it is often necessary to generate particle velocity distributions by using random numbers (random variates) at the initialization stage. It is straightforward to produce the Maxwell distribution, because one can employ a random number generator (RNG) for the normal (Gaussian) distribution, which is built-in in numerous programming languages and libraries. One can also implement the Box–Muller method for generating two independent normal variates from two uniform variates. Typically, three independent normal variates per particle are necessary to initialize a Maxwellian plasma. Related to this, one of the authors has developed an inverse sampling method for the 3D Maxwell distribution.(Umeda, 2024) The author approximated the cumulative distribution function of the Maxwellian by using an invertible function, and then constructed an inverse transform procedure.
@@
The standard Kappa distribution can be obtained by dividing three normal variates by a single gamma variate.(Abdul & Mace, 2015; Zenitani & Nakano, 2022) This is similar to an RNG for Student’s $t$-distribution, because the Kappa distribution is a multivariate $t$-distribution. It requires a gamma generator. In case it is unavailable, one can implement one of many gamma generators.(Devroye, 1986; Luengo, 2022) Alternatively, one can generate the standard Kappa distribution by using an acceptance–rejection method from the Type II Pareto distribution.(Zenitani, 2025) Importantly, most gamma generators internally employ the rejection method.
@@
Due to their excellent Watt-to-performance ratio (computational power per electric power consumption), high-performance computing on Graphics Processing Units (GPUs) has become increasingly important. Typical GPUs employ the Single-Instruction, Multiple-Threads (SIMT) execution model, in which a single instruction controls a group of threads simultaneously. This model is efficient at executing the same calculation in parallel; however, it is not good at handling loops whose iteration counts differ from thread to thread. This is the case for the rejection method. This may be one reason why numerical libraries on GPUs such as cuRAND and hipRAND do not provide gamma RNGs, even though they offer uniform and normal RNGs.
@@
Because of a growing demand for kinetic modeling and ongoing transition to GPU-based computers, we desire a GPU-friendly RNG for the Kappa distribution. In this study, we propose a novel Kappa generator, inspired by our previous work on the Maxwell distribution.(Umeda, 2024) Approximating the cumulative distribution function with an invertible function, we propose an inverse transform method. The algorithm is GPU-friendly, because it has no control flow divergence. It generates an approximate distribution that is indistinguishable from the Kappa distribution unless a very large number of particles are used.
@@
The remainder of this paper is organized as follows. Section II presents the standard Kappa distribution. Section III derives our approximate Kappa distribution, starting from the CDF of the Kappa distribution. Section IV constructs the inverse transform procedure. Section V evaluates the accuracy and performance of the new method by benchmarks. Section VI discusses the results with respect to the SIMT execution model, and then concludes this paper.
@@ II Kappa distribution
The phase space density of the Kappa distribution is
@@
where $N_{\kappa}$ is the number density, $\theta$ is the most probable speed, $\Gamma(x)$ is the gamma function, and $\kappa>3/2$ is the kappa index. Introducing a new variable $x\equiv v^{2}/\theta^{2}$ in the polar coordinates ($d^{3}v\rightarrow 4\pi v^{2}dv$), we translate the Kappa distribution into a beta-prime distribution, also known as the beta distribution of the second kind, with shape $(3/2,\kappa-1/2)$:
$f(x)dx$ $=\frac{N_{\kappa}}{\kappa\,B(3/2,\kappa-1/2)}\left(\frac{x}{\kappa}\right)^{1/2}\Big(1+\frac{x}{\kappa}\Big)^{-(\kappa+1)}dx$ (2)
Here, $B(x,y)$ is the beta function. Note that Eq. (2) contains a length scale of $\kappa$. The cumulative distribution function (CDF) of this distribution yields
@@
$=\frac{3\kappa}{2(2\kappa-3)}N_{\kappa}m\theta^{2}$
$~~~\times\left\{\frac{1}{\kappa B(5/2,\kappa-3/2)}\left(\frac{x}{\kappa}\right)^{3/2}\Big(1+\frac{x}{\kappa}\Big)^{-(\kappa+1)}\right\}\,dx$ (4)
@@
$\mathcal{E}_{\kappa}=\int_{0}^{\infty}f_{E}(x)dx=\frac{3\kappa}{2(2\kappa-3)}N_{\kappa}m\theta^{2}$ (5)
@@ III Approximate Kappa distribution
In line of Umeda (2024), we introduce the following approximation of the CDF:
@@
is the $q$-exponential function with the entropy index $q$. We set $q^{*}=1+1/\kappa^{*}$, because the Kappa distribution theory employs $q=1+1/\kappa$. Eq. (7) has four parameters of $a,b,c,$ and $q^{*}(\kappa^{*})$.
By differentiating $G(x)$, we obtain approximate distribution functions in the $x$ space and in the phase space:
$g(x)dx$ $=\dfrac{3N_{\kappa}(a+2bx+bcx^{2})}{2\left(1+cx\right)^{2}}$
$~~\times\sqrt{1-\left(1+\frac{ax+bx^{2}}{\kappa^{*}(1+cx)}\right)^{-\kappa^{*}}}$
$~~\times\left(1+\frac{ax+bx^{2}}{\kappa^{*}(1+cx)}\right)^{-(\kappa^{*}+1)}dx$ (9)
$g(\boldsymbol{v})d^{3}v$ $=\dfrac{3N_{\kappa}(a\theta^{4}+2bv^{2}\theta^{2}+bcv^{4})}{4\pi v\theta^{2}\left(\theta^{2}+cv^{2}\right)^{2}}$
$~~\times\sqrt{1-\left(1+\frac{(a\theta^{2}+bv^{2})v^{2}}{\kappa^{*}(\theta^{2}+cv^{2})\theta^{2}}\right)^{-\kappa^{*}}}$
$~~\times\left(1+\frac{(a\theta^{2}+bv^{2})v^{2}}{\kappa^{*}(\theta^{2}+cv^{2})\theta^{2}}\right)^{-(\kappa^{*}+1)}d^{3}v$ (10)
Figure 1: The two parameters, $a$ (Eq. (13)) and $b/c$ (Eq. (17)), are displayed as a function of $\kappa$.
Next, we seek for the four parameters. We consider Taylor expansions of Eqs. (3) and (10) near $x=0$:
$F(x)$ $=\frac{2}{3B(3/2,\kappa-1/2)}\left(\frac{x}{\kappa}\right)^{3/2}\left(1-\frac{\kappa+6}{5\kappa}x+\cdots\right)$ (11)
$G(x)$ $=a^{3/2}x^{3/2}\left\{1+\frac{3}{2}\left(\frac{b-ac}{a}-\frac{(2\kappa+1)a}{2(2\kappa-1)}\right)x+\cdots\right\}$ (12)
Comparing the coefficients of the $x^{3/2}$ terms, we find
$a=\frac{1}{\kappa}\left(\frac{2}{3B(3/2,\kappa-1/2)}\right)^{2/3}$ (13)
Also comparing Taylor expansions for $x\rightarrow\infty$,
$F(x)$ $=1-\frac{\kappa^{\kappa-1/2}}{(\kappa-1/2)B(3/2,\kappa-1/2)}x^{1/2-\kappa}+\cdots$ (14)
$G(x)$ $=1-\frac{3}{2}\left(\frac{(\kappa-1/2)c}{b}\right)^{\kappa^{*}}x^{-\kappa^{*}}+\cdots$ (15)
@@
$\frac{b}{c}$ $=\left(\kappa^{*}\frac{3}{2}B(3/2,\kappa-1/2)\right)^{1/\kappa^{*}}\frac{\kappa^{*}}{\kappa}$ (17)
The two parameters $a$ and $b/c$ are shown in Figure 1 as a function of $\kappa$. Starting from $2/3$ at $\kappa=3/2$, they increase monotonically, intersect each other at $\kappa\approx 4.1$, and then approach to their asymptotic limits. In the $\kappa\rightarrow\infty$ limit, considering ${\rm B}(\alpha,\beta)\approx\Gamma(\alpha)\beta^{-\alpha}$ for a large $\beta$, we find
@@
It approaches zero when the two distributions are similar, while it increases as the target probability distribution $g(x)$ deviates from the baseline probability distribution $f(x)$. We have carried out a survey of $\mathcal{D}$ in the 2D parameter space of $c\in[0,1.5]$ and $\kappa\in[1.5,10]$. Then, for a given $\kappa$, we look for $c$ that minimizes $\mathcal{D}$. Figure 2 shows $\mathcal{D}$ as a function of $(\kappa,c)$. The black curve shows an optimum $c$ that minimizes $\mathcal{D}$. The optimum $c$ diverges to infinity, before it returns to zero at $\kappa\approx 4.1$. This suggests that the fractional equation $(ax+bx^{2})/(1+cx)$ is approximated by $bx/c$, and then the equation is subsequently approximated by $ax$. Here, as shown in Figure 1, the two factors $a$ and $b/c$ are similar at $\kappa\approx 4.1$, and therefore it is less meaningful to choose $c$. Considering this, we look for a mathematical function that seamlessly connects the optimum $c$ for $\kappa\ll 4$ and for $\kappa\gg 4$. Using a symbolic regression library PySR(Cranmer, 2023), we find the following approximate formula from the optimum $c$ dataset in $\kappa\in[1.5,3.6]$ and $[4.6,10.0]$, and an artificial anchor point of $(\kappa,c)=(4.1,0.4)$.
@@
Figure 4(a) shows the relative entropy $\mathcal{D}$ by our approximation. This corresponds to the relative entropy along the approximate line (the red dashed line) in Figure 2, $\mathcal{D}(f\|g;\kappa,c_{\rm approx})$. For reference, the relative entropy along the optimum line (the black line in Fig. 2) is also presented by the dotted line. If we take a closer look, the optimum case looks somewhat jagged and it even surpasses the approximate case. These are due to the limited resolution of our grid search. Aside from these, the two entropies look similar, suggesting that the handy function (Eq. (20)) works excellently. The relative entropy is remarkably low for $\kappa<4$. This suggests that the two distributions are similar. It drops near $\kappa\approx 2.6$, but remains non-zero, because there is a subtle difference between the two distributions. The relative entropy has a peak at $\kappa\approx 4.2$. For $\kappa\gtrsim 5$, $\mathcal{D}$ drops to $\approx 10^{-6}$, and then it starts to increase very slowly. The three circles indicate values for $\kappa=3.0,4.1,$ and $7.5$, corresponding to the three representative cases in Figure 3(a). They effectively represent the three regimes of low-$\kappa$, high-$\kappa$, and the transition between them.
@@
We numerically calculate $\mathcal{E}_{\rm approx}$, and then we evaluate the difference to that of the Kappa distribution (Eq. (5)). Figure 4(b) shows the relative error in energy, $\pm{\Delta\mathcal{E}}/{\mathcal{E}_{\kappa}}=\pm({\mathcal{E}_{\rm approx}-\mathcal{E}_{\kappa}})/\mathcal{E}_{\kappa}$, and the positive and negative errors are shown by the solid and dashed lines. The error sometimes becomes positive, and sometimes negative. It becomes zero at $\kappa\approx 2.5$, which is slightly different from the minimum entropy point. The energy density matches that of the Kappa distribution within an error of $10^{-3}$, except for the largest error at $\kappa\approx 4.1$. Let us evaluate the anomaly at $\kappa\approx 4.1$ — from Figure 3(b), we see that plasma density is underestimated by 1% in the high-energy region of $v>3\theta$. This suggests 1% of the tail population in $v>3\theta$ is erroneously lost and redistributed in the low-energy region. From Eq. (6) with $x=3^{2}$ and $\kappa=4.1$, we see that the Kappa distribution carries $1-F_{\rm E}(x)\approx 19$% of the entire energy in this tail of $v>3\theta$. Since 1% of the tail population is lost and redistributed, one can estimate that 0.190.19% of the energy density is erroneously lost. This is in good agreement with the energy loss of $\mathcal{O}(10^{-2.5})$ near $\kappa\approx 4.1$ in Figure 4(b).

### v2

@@ Abstract
Kappa distribution, Particle-in-cell simulations, Numerical method, Monte Carlo method
@@ I Introduction
In space physics, the Kappa distribution is arguably the most important velocity distribution after the Maxwell distribution. Since it was empirically introduced during the 1960s (Olbert, 1968; Vasyliunas, 1968), the Kappa distribution has been extensively observed throughout the heliosphere (Pierrard & Lazar, 2010; Livadiotis, 2017) and it has become increasingly popular. Theoretically, it has also been established that the Kappa distribution is related to Tsallis statistics (Livadiotis & McComas, 2009). Since the Kappa distribution maximizes non-extensive entropy, it may represent a terminal state after plasma long-range interactions (Tsallis, 2023). Driven by spacecraft observations and various theoretical motivations, the kinetic modeling of Kappa-distributed plasmas is growing in its importance (Park et al., 2013; Gedalin & Ganushkina, 2022; Fitzmaurice, 2024; Nechaev et al., 2025).
@@
In kinetic plasma simulations, it is often necessary to generate particle velocity distributions by using random numbers (random variates) at the initialization stage. It is straightforward to produce the Maxwell distribution, because one can employ a random number generator (RNG) for the normal (Gaussian) distribution, which is built-in in numerous programming languages and libraries. One can also implement the Box–Muller method for generating two independent normal variates from two uniform variates. Typically, three independent normal variates per particle are necessary to initialize a Maxwellian plasma. Related to this, one of the authors has developed an inverse sampling method for the 3D Maxwell distribution (Umeda, 2024). The author approximated the cumulative distribution function of the Maxwellian by using an invertible function, and then constructed an inverse transform procedure.
@@
The standard Kappa distribution can be obtained by dividing three normal variates by a single gamma variate (Abdul & Mace, 2015; Zenitani & Nakano, 2022). This is similar to an RNG for Student’s $t$-distribution, because the Kappa distribution is a multivariate $t$-distribution. It requires a gamma generator. In case it is unavailable, one can implement one of many gamma generators (Devroye, 1986; Luengo, 2022). Alternatively, one can generate the standard Kappa distribution by using an acceptance–rejection method from the Type II Pareto distribution (Zenitani, 2025). Importantly, most gamma generators internally employ the rejection method.
@@
Due to their excellent performance per watt (computational power per electric power consumption), high-performance computing on Graphics Processing Units (GPUs) has become increasingly important. Typical GPUs employ the Single-Instruction, Multiple-Threads (SIMT) execution model, in which a single instruction controls a group of threads simultaneously. This model is efficient at executing the same calculation in parallel; however, it is not good at handling loops whose iteration counts differ from thread to thread. This is the case for the rejection method. This may be one reason why numerical libraries on GPUs such as cuRAND and hipRAND do not provide gamma RNGs, even though they offer uniform and normal RNGs.
@@
Because of a growing demand for kinetic modeling and ongoing transition to GPU-based computers, we desire a GPU-friendly RNG for the Kappa distribution. In this study, we propose a novel Kappa generator, inspired by our previous work on the Maxwell distribution (Umeda, 2024). Approximating the cumulative distribution function with an invertible function, we propose an inverse transform method. The algorithm is GPU-friendly, because it has no control flow divergence. It generates an approximate distribution that is indistinguishable from the Kappa distribution unless a very large number of particles are used.
@@
The remainder of this paper is organized as follows. Section 2 presents the standard Kappa distribution. Section 3 derives our approximate Kappa distribution, starting from the CDF of the Kappa distribution. Section 4 constructs the inverse transform procedure. Section 5 evaluates the accuracy and performance of the new method by benchmarks. Section 6 discusses the results with respect to the SIMT execution model, and then concludes this paper.
@@ II Kappa distribution
The phase space density of the (nonrelativistic) Kappa distribution is
@@
where $N_{\kappa}$ is the number density, $\theta$ is the most probable speed, $\Gamma(x)$ is the gamma function, and $\kappa>3/2$ is the kappa index. In this study, we discuss the Kappa distribution as a probability distribution function; thus, we hereafter set $N_{\kappa}=1$. Introducing a new variable $x\equiv v^{2}/\theta^{2}$ in the polar coordinates ($d^{3}v\rightarrow 4\pi v^{2}dv$), we translate the Kappa distribution into a beta-prime distribution, also known as the beta distribution of the second kind, with shape $(3/2,\kappa-1/2)$:
$f(x)dx$ $=\frac{1}{\kappa\,B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(\frac{x}{\kappa}\right)^{1/2}\Big(1+\frac{x}{\kappa}\Big)^{-(\kappa+1)}dx$ (2)
Here, $B(x,y)$ is the beta function. Note that Eq. (2) contains a scaling factor $\kappa$. The cumulative distribution function (CDF) of this distribution yields
@@
$=\frac{3\kappa}{2(2\kappa-3)}m\theta^{2}\left\{\frac{1}{\kappa B\left(\frac{5}{2},\kappa-\frac{3}{2}\right)}\left(\frac{x}{\kappa}\right)^{3/2}\Big(1+\frac{x}{\kappa}\Big)^{-(\kappa+1)}\right\}\,dx$ (4)
@@
$\mathcal{E}_{\kappa}=\int_{0}^{\infty}f_{E}(x)dx=\frac{3\kappa}{2(2\kappa-3)}m\theta^{2}$ (5)
@@ III Approximate Kappa distribution
In line with Umeda (2024), we introduce the following approximation of the CDF:
@@
is the $q$-exponential function with the entropy index $q$. We use fitting parameters $q^{*}\equiv 1+1/\kappa^{*}$ and $\kappa^{*}$, because the entropy index is expected to be $q=1+1/\kappa$ in Kappa distribution theory. Eq. (7) has four parameters: $a,b,c,$ and $\kappa^{*}(q^{*})$.
By differentiating $G(x)$, we obtain approximate distribution functions in the $x$ space and in the velocity space.
$g(x)dx$ $=\dfrac{3(a+2bx+bcx^{2})}{2\left(1+cx\right)^{2}}\sqrt{1-\left(1+\frac{ax+bx^{2}}{\kappa^{*}(1+cx)}\right)^{-\kappa^{*}}}\left(1+\frac{ax+bx^{2}}{\kappa^{*}(1+cx)}\right)^{-(\kappa^{*}+1)}dx$ (9)
$g(\boldsymbol{v})d^{3}v$ $=\dfrac{3(a\theta^{4}+2bv^{2}\theta^{2}+bcv^{4})}{4\pi v\theta^{2}\left(\theta^{2}+cv^{2}\right)^{2}}\sqrt{1-\left(1+\frac{(a\theta^{2}+bv^{2})v^{2}}{\kappa^{*}(\theta^{2}+cv^{2})\theta^{2}}\right)^{-\kappa^{*}}}\left(1+\frac{(a\theta^{2}+bv^{2})v^{2}}{\kappa^{*}(\theta^{2}+cv^{2})\theta^{2}}\right)^{-(\kappa^{*}+1)}d^{3}v$ (10)
Figure 1: The two parameters, $a$ (Eq. (13)) and $b/c$ (Eq. (17)), are displayed as functions of $\kappa$.
Next, we seek the four parameters. We consider Taylor expansions of Eqs. (3) and (7) near $x=0$.
$F(x)$ $=\frac{2}{3B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\left(\frac{x}{\kappa}\right)^{3/2}\left(1-\frac{3(\kappa+1)}{5\kappa}x+\mathcal{O}(x^{2})\right)$ (11)
$G(x)$ $=a^{3/2}x^{3/2}\left\{1+\frac{3}{2}\left(\frac{b-ac}{a}-\frac{(\kappa^{*}+1)a}{2\kappa^{*}}\right)x+\mathcal{O}(x^{2})\right\}$ (12)
Their derivations are given in the Appendix. Comparing the coefficients of the $x^{3/2}$ terms, we find
$a=\frac{1}{\kappa}\left(\frac{2}{3B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}\right)^{2/3}$ (13)
Similarly comparing Taylor expansions for $x\rightarrow\infty$ (also derived in the Appendix),
$F(x)$ $=1-\frac{\kappa^{\kappa-1/2}}{(\kappa-1/2)B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)}x^{-(\kappa-1/2)}+\mathcal{O}(x^{-(\kappa+1/2)})$ (14)
$G(x)$ $=1-\frac{3}{2}\left(\frac{c\kappa^{*}}{b}\right)^{\kappa^{*}}x^{-\kappa^{*}}+\mathcal{O}(x^{-(\kappa^{*}+1)})$ (15)
@@
$\frac{b}{c}$ $=\left\{\kappa^{*}\frac{3}{2}B\left(\frac{3}{2},\kappa-\frac{1}{2}\right)\right\}^{1/\kappa^{*}}\frac{\kappa^{*}}{\kappa}$ (17)
The two parameters $a$ and $b/c$ are shown in Figure 1 as functions of $\kappa$. Starting from $2/3$ at $\kappa=3/2$, they increase monotonically, intersect each other at $\kappa\approx 4.1$, and then approach to their asymptotic limits. In the $\kappa\rightarrow\infty$ limit, considering $B(\alpha,\beta)\approx\Gamma(\alpha)\beta^{-\alpha}$ for a large $\beta$, we find
@@
It approaches zero when the two distributions are similar, while it increases when the target probability distribution $g(x)$ deviates from the baseline probability distribution $f(x)$. We have carried out a survey of $\mathcal{D}$ in the 2D parameter space of $c\in[0,1.5]$ and $\kappa\in[1.5,10]$. Then, for a given $\kappa$, we look for $c$ that minimizes $\mathcal{D}$. Figure 2 shows $\mathcal{D}$ as a function of $(\kappa,c)$. The black curve shows an optimum $c$ that minimizes $\mathcal{D}$. The optimum $c$ diverges to infinity, before it returns to zero at $\kappa\approx 4.1$. This suggests that an optimum value of $(ax+bx^{2})/(1+cx)$ is $\approx bx/c$ in the left vicinity of $\kappa\approx 4.1$ and $\approx ax$ in the right vicinity. This fractional term is continuous, because the two factors $a$ meets $b/c$ at $\kappa\approx 4.1$ (Fig. 1). The parameter $c$ essentially controls the weights of $bx/c$ and $ax$ inside the fraction equation. When the two factors $a$ and $b/c$ are similar, it is less meaningful to choose $c$ near $\kappa\approx 4.1$. Considering this, we look for a mathematical function that seamlessly connects the optimum $c$ for $\kappa\ll 4$ and for $\kappa\gg 4$. Using a symbolic regression library PySR (Cranmer, 2023), we find the following approximate formula from the optimum $c$ dataset in $\kappa\in[1.5,3.6]$ and $[4.6,10.0]$, and an artificial anchor point of $(\kappa,c)=(4.1,0.4)$.
@@
Figure 4(a) shows the relative entropy $\mathcal{D}$ of our approximation. This corresponds to the relative entropy along the approximate line (the red dashed line) in Figure 2, $\mathcal{D}(f\|g;\kappa,c_{\rm approx})$. For reference, the relative entropy along the optimum line (the black line in Fig. 2) is also presented by the gray dashed line. If we take a closer look, the optimum case looks somewhat jagged and it even surpasses the approximate case. These are due to the limited resolution of our grid search. Aside from these, the two entropies look similar, suggesting that the handy function (Eq. (20)) works excellently. The relative entropy is remarkably low for $\kappa<4$. This suggests that the two distributions are similar. It drops near $\kappa\approx 2.6$, but remains non-zero, because there is a subtle difference between the two distributions. The relative entropy has a peak at $\kappa\approx 4.2$. For $\kappa\gtrsim 5$, $\mathcal{D}$ drops to $\approx 10^{-6}$, and then it starts to increase very slowly. The three circles indicate values for $\kappa=3.0,4.1,$ and $7.5$, corresponding to the three representative cases in Figure 3(a). They effectively represent the three regimes of low-$\kappa$, high-$\kappa$, and the transition between them.
@@
We numerically calculate $\mathcal{E}_{\rm approx}$, and then we evaluate the difference to that of the Kappa distribution (Eq. (5)). Figure 4(b) shows the relative error in energy, $\pm{\Delta\mathcal{E}}/{\mathcal{E}_{\kappa}}=\pm({\mathcal{E}_{\rm approx}-\mathcal{E}_{\kappa}})/\mathcal{E}_{\kappa}$, and the positive and negative errors are shown by the solid and dashed lines. The error sometimes becomes positive, and sometimes negative. It becomes zero at $\kappa\approx 2.5$, which is slightly different from the minimum entropy point. The energy density matches that of the Kappa distribution within an error of $10^{-3}$, except for the largest error at $\kappa\approx 4.1$. Let us evaluate the anomaly at $\kappa\approx 4.1$ — from Figure 3(b), we see that plasma density is underestimated by 1% in the high-energy region of $v>3\theta$. This suggests that 1% of the tail population in $v>3\theta$ is erroneously lost and redistributed in the low-energy region. From Eq. (6) with $x=3^{2}$ and $\kappa=4.1$, we see that the Kappa distribution carries $1-F_{\rm E}(x)\approx 19$% of the entire energy in this tail of $v>3\theta$. Since 1% of the tail population is lost and redistributed, one can estimate that 0.190.19% of the energy density is erroneously lost. This is in good agreement with the energy loss of $\mathcal{O}(10^{-2.5})$ near $\kappa\approx 4.1$ in Figure 4(b).